The moon Titan orbits the planet Saturn with a period of 1.9 × 106 s. The average radius of

this orbit is 1.6 × 109 m.

a) What is Titan’s centripetal acceleration?

b) What is Titans Velocity?


1
Expert's answer
2021-06-09T08:05:06-0400

The information given is

The period of a complete orbit is "T=1.9\times 10^{6}\;\text{s}"

The average radius of the orbit is. "r=1.6\times 10^{9}\;\text{s}"


Considering the circular orbit.


The angular speed is given by

"w=\dfrac{2\;\pi\;\text{rad}}{T}"


Where.

"2\;\pi\;\text{rad}" is the angular displacement of a complete orbit

"T" is the period of a complete orbit.


Evaluating numerically.

"w=\dfrac{2\;\pi\;\text{rad}}{T}\\
w=\dfrac{2\;\pi\;\text{rad}}{1.9\times 10^{6}\;\text{s} }\\
w= 3.3\times 10^{-6} \text{rad}/\text{s}"



The centripetal acceleration is given by.

"a_{c}=w^{2}\;r"


Where.

"w" is the angular speed

"r" is the radius.


Evaluating numerically.

"a_{c}=w^{2}\;r\\
a_{c}=( 3.3\times 10^{-6} \text{rad}/\text{s} )^{2}\times 1.6\times 10^{9}\;\text{s} \\
a_{c}=1.7\times 10^{-2}\;\text{m}/\text{s}^{2}"


Answer A

The centripetal acceleration is "\displaystyle \color{red}{\boxed{a_{c}=1.7\times 10^{-2}\;\text{m}/\text{s}^{2}}}"


Part b


The linear velocity is given by

"V=w\cdot r"


Where


"w" is the angular speed

"r" is the radio


Evaluating numerically

"V=w\cdot r\\
V= 3.3\times 10^{-6} \text{rad}/\text{s}\times 1.6\times 10^{9}\;\text{s}\\
V=5.3\times 10^{3}\;\text{m}/\text{s}"


Answer B

The velocity is "\displaystyle \color{red}{\boxed{V=5.3\times 10^{3}\;\text{m}/\text{s}}}"